# Minimal-doubling and single-Weyl Hamiltonians

### v1

@@ Abstract
We develop a systematic Hamiltonian formulation of minimally doubled lattice fermions in $(3+1)$ dimensions, derive their nodal structures (structures of zeros), and classify their symmetry patterns for both four-component Dirac and two-component Weyl constructions. Motivated by recent single-Weyl proposals based on Bogoliubov-de Gennes (BdG) representation, we argue that the corresponding single-Weyl Hamiltonians are obtained from the minimal-doubling Hamiltonians supplemented by an appropriate species-splitting mass term, and we re-examine the non-onsite symmetry protecting the physical Weyl node in terms of a Ginsparg–Wilson–type relation. We then construct a one-parameter family of deformations that preserves all the symmetries and demonstrate that, once the parameter exceeds a critical value, additional Weyl nodes emerge and the system exits the single-node regime. This indicates that in interacting theories radiative corrections can generate symmetry-allowed counterterms, so maintaining the desired single-Weyl phase generically requires “moderate” parameter tuning.
@@ I Introduction
Formulating chiral fermions on a lattice is a long-standing problem at the intersection of quantum field theory, condensed-matter band theory, and nonperturbative gauge dynamics Wilson:1974sk ; Creutz:1980zw . The Nielsen–Ninomiya no-go theorem implies that, under standard assumptions such as onsite chiral symmetry, locality, Hermiticity, and translation invariance, lattice fermions necessarily appear in left–right pairs, forbidding an isolated Weyl mode Karsten:1980wd ; Nielsen:1980rz ; Nielsen:1981xu ; Nielsen:1981hk . Well-established lattice regularizations of vector-like gauge theories, such as Wilson fermions and staggered fermions, achieve practical control of doublers at the cost of either explicitly breaking chiral symmetry or introducing additional tastes Wilson:1975id ; Kogut:1974ag ; Susskind:1976jm ; Kawamoto:1981hw ; Sharatchandra:1981si ; Golterman:1984cy ; Golterman:1985dz ; Kilcup:1986dg . Domain-wall and overlap fermions realize single-flavor setups at the cost of breaking locality or onsite chiral symmetry Kaplan:1992bt ; Shamir:1993zy ; Furman:1994ky ; Neuberger:1998wv ; Ginsparg:1981bj ; Luscher:1998pqa . The improved and combined approaches of Wilson and staggered fermions have also been proposed, including the generalized Wilson fermions Bietenholz:1999km ; Creutz:2010bm ; Durr:2010ch ; Durr:2012dw ; Misumi:2012eh ; Cho:2013yha ; Cho:2015ffa ; Durr:2017wfi , the staggered-Wilson fermions Adams:2009eb ; Adams:2010gx ; Hoelbling:2010jw ; deForcrand:2011ak ; Creutz:2011cd ; Misumi:2011su ; Follana:2011kh ; deForcrand:2012bm ; Misumi:2012sp ; Misumi:2012eh ; Durr:2013gp ; Hoelbling:2016qfv ; Zielinski:2017pko , and the central-branch Wilson fermion Kimura:2011ik ; Creutz:2011cd ; Misumi:2012eh ; Chowdhury:2013ux ; Misumi:2019jrt ; Misumi:2020eyx . By contrast, minimally doubled fermions realize the smallest allowed number of species within the no-go theorem while preserving an exact onsite chiral symmetry, typically by sacrificing full hypercubic invariance and other discrete symmetries Karsten:1981gd ; Wilczek:1987kw ; Creutz:2007af ; Borici:2007kz ; Bedaque:2008xs ; Bedaque:2008jm ; Capitani:2009yn ; Kimura:2009qe ; Kimura:2009di ; Creutz:2010cz ; Creutz:2010qm ; Capitani:2010nn ; Tiburzi:2010bm ; Kamata:2011jn ; Misumi:2012uu ; Misumi:2012ky ; Capitani:2013zta ; Capitani:2013iha ; Misumi:2013maa ; Weber:2013tfa ; Weber:2017eds ; Durr:2020yqa .
@@
Another motivation for the present work arises from recent proposals aiming to isolate a single Weyl node by enlarging the Hilbert space through a Bogoliubov–de Gennes (BdG) or Nambu particle/hole doubling and adding a momentum-dependent pairing term in the Hamiltonian formalism Gioia:2025bhl . Such models can possess exact, momentum-dependent conserved charges that commute with the lattice Hamiltonian while acting as a non-onsite generator in the position space. From the viewpoint of lattice chiral symmetry, these protecting charges are naturally interpreted as Hamiltonian analogues of Ginsparg–Wilson-type chiral structures Creutz:2001wp ; Fujikawa:2000my ; Clancy:2023kla ; Clancy:2023ino ; Fidkowski:2023sif ; Singh:2025sye : they are conserved, but their spectra need not be quantized, and the associated “chiral” transformations are generically momentum- or position-dependent. Classification of the similar single-Weyl Hamiltonians are found in Ref. Meyniel:2025euu . For the attempt to define chiral fermions in staggered formulations and the related works, see Chatterjee:2024gje ; Onogi:2025xir ; Aoki:2025vtp . (See also Ref. Eichten:1985ft for the old related work.)
The first aim of this work is therefore methodological: we provide a unified Hamiltonian construction of minimal-doubling fermions and systematically identify which discrete symmetries are preserved or broken in the resulting Hamiltonians. The second aim is to clarify the relation between the minimal-doubling Hamiltonian and the recently proposed BdG-based single-Weyl Hamiltonians, and to analyze the latter in terms of their chiral structure and possible symmetry-allowed deformations. From the lattice-gauge-theory perspective, the key question is whether the single-node phase can be maintained in the interacting theory.
@@ II.1 Karsten-Wilczek type
For exact chiral symmetry, the one-particle Hamiltonian $h$ (Bloch Hamiltonian) must commute with $\gamma_{5}$ as $h,\gamma_{5}]=0$. Any term proportional to $\beta=\gamma_{0}$ anticommutes with $\gamma_{5}$ and is forbidden. Thus, the Hamiltonian must be a linear combination of the $\alpha_{j}=\gamma_{0}\gamma_{j}$ $(j=1,2,3)$ matrices only.
II.1.1 Dirac fermion
We first consider a cubic spatial lattice with sites ${\bm{x}}=(x,y,z)\in\mathbb{Z}^{3}$ and a four-component Dirac spinor $\psi_{{\bm{x}}}$ at each site with the periodic boundary condition. Gamma matrices $\gamma_{\mu}$ or $\alpha_{j}=\gamma_{0}\gamma_{j}$ act on the Dirac spinor. The position-space and momentum-space representations of the Dirac spinor have the relation
@@ II.1.1 Dirac fermion
Therefore, from the condensed-matter viewpoint, this model describes a phase analogous to a PT-symmetric Dirac semimetal. Although the individual P and T symmetries are broken, the combined PT symmetry, along with the chiral symmetry, ensures that the bands remain doubly degenerate throughout the Brillouin zone. In addition to the standard discrete symmetries, one can define the special parity-like symmetry, which can be regarded as a Hamiltonian version of mirror fermion symmetry discussed for the KW Dirac operator Weber:2017eds
@@
II.1.2 Weyl fermion
We now consider a two-component Weyl spinor $\chi_{{\bm{x}}}$ at each site. Pauli matrices $\sigma_{1,2,3}$ act on the Weyl spinor. The position-space and momentum-space representations are related as $\chi_{\bm{x}}=\int_{-\pi}^{\pi}\frac{d^{3}p}{(2\pi)^{3}}\,e^{i{\bm{p}}\cdot{\bm{x}}}\chi_{\bm{p}}$. The massless and free KW Hamiltonian of the 2-component Weyl spinor is
@@ II.1.2 Weyl fermion
This Hamiltonian breaks the cubic rotational symmetry to the two-dimensional subgroup, and also breaks both Parity and Time-reversal invariances. In this sense, this Hamiltonian corresponds to the P-broken magnetic Weyl semimetal in the condensed-matter context. In comparison to the case of the Dirac formalism, the Weyl KW formalism does not possess the combined PT invariance, while it still has the special parity-like symmetry $\tilde{\rm P}:\,{\bm{p}}\mapsto(-p_{1},-p_{2},-p_{3}+\pi),\,\chi\mapsto\sigma_{3}\,\chi$.
@@
As shown in Appendix A, one can introduce the species-splitting mass term to the minimal-doubling fermion based on the flavor representation of point-split fields Creutz:2010bm . Such a mass term breaks the $U(1)$ onsite symmetry and gaps one of the two Weyl nodes with keeping the other massless. For instance, by introducing the species-splitting mass term into the KW Weyl Hamiltonian (19), one reaches a single-Weyl Hamiltonian
@@
which keeps the Weyl node at ${\bm{p}}_{A}=(0,0,0)$ massless while gaps the other at ${\bm{p}}_{B}=(0,0,\pi)$. In this setup, however, the $U(1)$ symmetry to protect the gapless node is absent and the mode at ${\bm{p}}_{A}=(0,0,0)$ can be gapped due to the radiative correction in the interacting theory. Thus, it is not a chiral Weyl fermion. In the chiral-symmetric Weyl setup, which will be discussed in Sec. III, the similar species-splitting mass term will be introduced not in the flavor representation but in the Bogoliubov–de Gennes (BdG) representation on the Nambu (particle/hole) space.
@@ II.2 Twisted-ordering type
It has two Weyl nodes at ${\bm{p}}_{A}=(0,0,0)$ and ${\bm{p}}_{B}=(\pi/2,\pi/2,\pi/2)$, yielding two gapless (massless) Weyl modes. While it breaks P and T symmetries, it possesses the $U(1)$ onsite symmetry, which corresponds to $U(1)$ vector symmetry of the single Dirac fermion composed of the two Weyl modes at ${\bm{p}}_{A},{\bm{p}}_{B}$ as in the KW Weyl Hamiltonian in Sec. II.1.2.
@@ II.3 Borici-Creutz type
with $\sigma^{\prime}_{j}=A_{ji}\sigma_{i}$. It has two Weyl nodes at ${\bm{p}}=(\pi/4,\pi/4,\pi/4)$ and ${\bm{p}}=-(\pi/4,\pi/4,\pi/4)$, giving two massless Weyl modes with the $U(1)$ onsite symmetry preserved and P, T symmetries broken.
@@ II.4 Classification and Generalization
where $\alpha_{j}^{\prime}=A_{ji}\alpha_{i}$. Here $\beta_{j}$ and $\alpha$ are related with $A$ as $\alpha=\sum_{j}\alpha_{j}\sin 2\beta_{j}=\sum_{j}\alpha_{j}^{\prime}\sin 2\beta_{j}$, which means that $\sin 2\beta_{j}$ is an eigenvector of $A$. Thus, once $A$ is fixed, $\beta_{j}$ and $\alpha$ are determined up to an overall factor. By imposing these conditions on $A$, $\beta_{j}$ and $\alpha$, the Hamiltonian can be minimal-doubling, where $\pm\beta_{j}$ correspond to locations of two nodes. The Hamiltonian (35) reduces to Borici-Creutz one (28) by choosing (29) for $A$, which fixes $\beta_{j}=\pi/4$ and $\alpha=\sum_{j=1}^{3}\alpha_{j}$. The Hamiltonian (35) also yields the twisted-ordering one by choosing $A$ as
@@ III.1 Review of Gioia–Thorngren setup
$h^{\text{\tiny BdG}}_{\text{\tiny KW}}({\bm{p}})\,=\,\frac{1}{2}\Bigl[\sigma_{1}\,\sin p_{1}+\sigma_{3}\,\sin k_{2}+\tau_{3}\sigma_{2}\,\sin k_{3}+\sigma_{2}\,r\left(2-\cos p_{1}-\cos p_{2}\right)\Bigr],$ (41)
@@ III.3 One-parameter deformation, extra nodes and need for tuning
Because $\sigma_{1},\sigma_{2},\sigma_{3}$ anticommute and the $\tau$ matrices commute with the $\sigma$ matrices, the spectrum follows from
$4h(\bm{p})^{2}$ $=\sin^{2}p_{1}+\sin^{2}p_{2}+\Big[rM\,\mathbf{1}_{\tau}+a\tau_{3}+b\tau_{1}\Big]^{2}$
$=\sin^{2}p_{1}+\sin^{2}p_{2}+\Big(rM\pm\sqrt{a^{2}+b^{2}}\Big)^{2},$ (59)
@@
Thus, once the bound $|\mu|=\sqrt{r^{2}-1}$ is violated, the system is no longer in the phase with a single Weyl node: additional gapless nodes are created at $(\pi,0,\hat{p}_{3})$ and $(0,\pi,\hat{p}_{3})$.

### v2

@@ Abstract
We develop a systematic Hamiltonian formulation of minimally doubled lattice fermions in $(3+1)$ dimensions, derive their nodal structures (structures of zeros), and classify their symmetry patterns for both four-component Dirac and two-component Weyl constructions. Motivated by recent single-Weyl proposals based on Bogoliubov-de Gennes (BdG) representation, we argue that the corresponding single-Weyl Hamiltonians are obtained from the minimal-doubling Hamiltonians supplemented by an appropriate species-splitting mass term, and we re-examine the non-onsite symmetry protecting the physical Weyl node in terms of a Ginsparg-Wilson-type relation. We then construct a one-parameter family of deformations that preserves all the symmetries and demonstrate that, once the parameter exceeds a critical value, additional Weyl nodes emerge and the system exits the single-node regime. This indicates that in interacting theories radiative corrections can generate symmetry-allowed counterterms, so maintaining the desired single-Weyl phase generically requires “moderate” parameter tuning.
@@ I Introduction
Formulating chiral fermions on a lattice is a long-standing problem at the intersection of quantum field theory, condensed-matter band theory, and nonperturbative gauge dynamics Wilson:1974sk ; Creutz:1980zw . The Nielsen-Ninomiya no-go theorem is most generally formulated in terms of conserved onsite charges with discrete spectra, together with locality, Hermiticity, and translation invariance. Under these standard assumptions, left- and right-handed gapless fermions must appear in matching numbers within each charge sector, thereby forbidding an isolated Weyl mode Karsten:1980wd ; Nielsen:1980rz ; Nielsen:1981xu ; Nielsen:1981hk . Well-established lattice regularizations of vector-like gauge theories, such as Wilson fermions and staggered fermions, achieve practical control of doublers at the cost of either explicitly breaking chiral symmetry (breaking a conserved onsite charge) or introducing additional tastes Wilson:1975id ; Kogut:1974ag ; Susskind:1976jm ; Kawamoto:1981hw ; Sharatchandra:1981si ; Golterman:1984cy ; Golterman:1985dz ; Kilcup:1986dg . Domain-wall and overlap fermions realize single-flavor setups at the cost of breaking locality or onsite chiral symmetry Kaplan:1992bt ; Shamir:1993zy ; Furman:1994ky ; Neuberger:1998wv ; Ginsparg:1981bj ; Luscher:1998pqa . The improved and combined approaches of Wilson and staggered fermions have also been proposed, including the generalized Wilson fermions Bietenholz:1999km ; Creutz:2010bm ; Durr:2010ch ; Durr:2012dw ; Misumi:2012eh ; Cho:2013yha ; Cho:2015ffa ; Durr:2017wfi , the staggered-Wilson fermions Adams:2009eb ; Adams:2010gx ; Hoelbling:2010jw ; deForcrand:2011ak ; Creutz:2011cd ; Misumi:2011su ; Follana:2011kh ; deForcrand:2012bm ; Misumi:2012sp ; Misumi:2012eh ; Durr:2013gp ; Hoelbling:2016qfv ; Zielinski:2017pko , and the central-branch Wilson fermion Kimura:2011ik ; Creutz:2011cd ; Misumi:2012eh ; Chowdhury:2013ux ; Misumi:2019jrt ; Misumi:2020eyx . By contrast, minimally doubled fermions realize the smallest allowed number of species within the no-go theorem while preserving an exact onsite chiral symmetry, typically by sacrificing full hypercubic invariance and other discrete symmetries Karsten:1981gd ; Wilczek:1987kw ; Creutz:2007af ; Borici:2007kz ; Bedaque:2008xs ; Bedaque:2008jm ; Capitani:2009yn ; Kimura:2009qe ; Kimura:2009di ; Creutz:2010cz ; Creutz:2010qm ; Capitani:2010nn ; Tiburzi:2010bm ; Kamata:2011jn ; Misumi:2012uu ; Misumi:2012ky ; Capitani:2013zta ; Capitani:2013iha ; Misumi:2013maa ; Weber:2013tfa ; Weber:2017eds ; Durr:2020yqa .
@@
Another motivation for the present work arises from recent proposals aiming to isolate a single Weyl node by rewriting the Hamiltonian in a doubled Bogoliubov-de Gennes (BdG) or Nambu particle/hole representation and adding a momentum-dependent pairing term in the Hamiltonian formalism Gioia:2025bhl . Such models can possess exact, momentum-dependent conserved charges that commute with the lattice Hamiltonian while acting as a non-onsite generator in the position space. From the viewpoint of lattice chiral symmetry, these conserved non-onsite charges are naturally interpreted as Hamiltonian analogues of Ginsparg–Wilson-type chiral symmetries Creutz:2001wp ; Fujikawa:2000my ; Clancy:2023kla ; Clancy:2023ino ; Fidkowski:2023sif ; Singh:2025sye : they are conserved, but their spectra need not be quantized, and the associated “chiral” transformations are generically momentum- or position-dependent. Classification of the similar single-Weyl Hamiltonians are found in Ref. Meyniel:2025euu . For the attempt to define chiral fermions in staggered formulations and the related works, see Chatterjee:2024gje ; Onogi:2025xir ; Aoki:2025vtp . (See also Ref. Eichten:1985ft for the old related work.)
The first aim of this work is therefore methodological: we provide a unified Hamiltonian construction of minimal-doubling fermions and systematically identify which discrete symmetries are preserved or broken in the resulting Hamiltonians. The second aim is to clarify the relation between the minimal-doubling Hamiltonian and the recently proposed BdG-based single-Weyl Hamiltonians, and to analyze the latter in terms of their chiral structure and possible symmetry-allowed deformations. From the lattice-gauge-theory perspective, the key question is whether the single-node phase can be maintained in the interacting theory. Throughout this paper, we use the term “node” to mean a gapless point (equivalently, a zero of the Bloch Hamiltonian / energy spectrum) in the Brillouin zone.
@@ II.1 Karsten-Wilczek type
For exact chiral symmetry, the one-particle Hamiltonian $h$ (Bloch Hamiltonian) must commute with $\gamma_{5}$ as $h,\gamma_{5}]=0$. Any term proportional to $\beta=\gamma_{0}$ anticommutes with $\gamma_{5}$ and is forbidden. In the restricted class considered here, namely, Hamiltonians continuously connected to the standard massless Dirac Hamiltonian, we restrict ourselves to the $\alpha_{j}=\gamma_{0}\gamma_{j}$ $(j=1,2,3)$ structures.
II.1.1 Four-component fermion
We first consider a cubic spatial lattice with sites ${\bm{x}}=(x,y,z)\in\mathbb{Z}^{3}$ and a four-component spinor $\psi_{{\bm{x}}}$ at each site with the periodic boundary condition. Gamma matrices $\gamma_{\mu}$ or $\alpha_{j}=\gamma_{0}\gamma_{j}$ act on the Dirac spinor. The position-space and momentum-space representations of the Dirac spinor have the relation
@@ II.1.1 Four-component fermion
Therefore, from the condensed-matter viewpoint, this model describes a phase analogous to a PT-symmetric Dirac semimetal. Since $h^{\text{\tiny D}}_{\text{\tiny KW}}$ is a $4\times 4$ Hamiltonian built from three mutually anti-commuting $\alpha$-matrices, its spectrum is $\pm E(p)$ with each energy level being two-fold degenerate. The PT symmetry is consistent with this structure, while the chiral symmetry gives the $E\leftrightarrow-E$ spectral symmetry. In addition to the standard discrete symmetries, one can define the special parity-like symmetry, which can be regarded as a Hamiltonian version of mirror fermion symmetry discussed for the KW Dirac operator Weber:2017eds
@@
II.1.2 Two-component fermion
We now consider a two-component spinor $\chi_{{\bm{x}}}$ at each site. Pauli matrices $\sigma_{1,2,3}$ act on the Weyl spinor. The position-space and momentum-space representations are related as $\chi_{\bm{x}}=\int_{-\pi}^{\pi}\frac{d^{3}p}{(2\pi)^{3}}\,e^{i{\bm{p}}\cdot{\bm{x}}}\chi_{\bm{p}}$. The massless and free KW Hamiltonian of the 2-component Weyl spinor is
@@ II.1.2 Two-component fermion
Here ${\bm{1}}_{\sigma}$ denotes the $2\times 2$ identity matrix acting on the Weyl-spinor space. This Hamiltonian breaks the cubic rotational symmetry to the two-dimensional subgroup, and also breaks both Parity and Time-reversal invariances. In this sense, this Hamiltonian corresponds to the P-broken magnetic Weyl semimetal in the condensed-matter context. In comparison to the case of the Dirac formalism, the Weyl KW formalism does not possess the combined PT invariance, while it still has the special parity-like symmetry $\tilde{\rm P}:\,{\bm{p}}\mapsto(-p_{1},-p_{2},-p_{3}+\pi),\,\chi\mapsto\sigma_{3}\,\chi$.
@@
As shown in Appendix A, one can introduce the species-splitting mass term to the minimal-doubling fermion based on the flavor representation of point-split fields Creutz:2010bm . Such a term gaps one of the two Weyl nodes while preserving the ordinary onsite fermion-number $U(1)$ symmetry. For instance, by introducing the species-splitting mass term into the KW two-component Hamiltonian (19), one reaches a single-Weyl Hamiltonian
@@
which keeps the Weyl node at ${\bm{p}}_{A}=(0,0,0)$ massless while gaps the other at ${\bm{p}}_{B}=(0,0,\pi)$. In this setup, however, the onsite fermion-number $U(1)$ symmetry does not work to protect the remaining gapless node. Thus, although Eq. (25) yields a single gapless Weyl node in the free theory, it can be gapped due to the radiative correction in the interacting theory, thus it does not provide a symmetry-protected chiral single-Weyl construction. In the chiral-symmetric Weyl setup discussed in Sec. III, the analogous species-splitting term is introduced not in flavor space but in the Bogoliubov-de Gennes (BdG) representation on the Nambu (particle/hole) space, where the onsite $U(1)$ is explicitly broken.
@@ II.2 Twisted-ordering type
It has two Weyl nodes at ${\bm{p}}_{A}=(0,0,0)$ and ${\bm{p}}_{B}=(\pi/2,\pi/2,\pi/2)$, yielding two gapless (massless) Weyl modes. They carry opposite physical chiralities (helicities), and therefore can be combined into a single massless Dirac fermion in the low-energy theory. While it breaks P and T symmetries, it possesses the $U(1)$ onsite symmetry, which corresponds to $U(1)$ vector symmetry of the single Dirac fermion composed of the two Weyl modes at ${\bm{p}}_{A},{\bm{p}}_{B}$ as in the KW Weyl Hamiltonian in Sec. II.1.2.
@@ II.3 Borici-Creutz type
with $\sigma^{\prime}_{j}=A_{ji}\sigma_{i}$. It has two Weyl nodes at ${\bm{p}}=(\pi/4,\pi/4,\pi/4)$ and ${\bm{p}}=-(\pi/4,\pi/4,\pi/4)$, giving two massless Weyl modes with the $U(1)$ onsite symmetry preserved and P, T symmetries broken. Again, the two Weyl modes carry opposite physical chiralities, and therefore can be combined into a single massless Dirac fermion in the low-energy theory.
@@ II.4 Classification and Generalization
where $\alpha_{j}^{\prime}=A_{ji}\alpha_{i}$. Here $\beta_{j}$ and $\alpha$ are related with AA as $\alpha=\sum_{j}\alpha_{j}\sin 2\beta_{j}=\sum_{j}\alpha_{j}^{\prime}\sin 2\beta_{j}$, which means that $\sin 2\beta_{j}$ is an eigenvector of AA. Thus, once AA is fixed, $\beta_{j}$ and $\alpha$ are determined up to an overall factor. By imposing these conditions on AA, $\beta_{j}$ and $\alpha$, the Hamiltonian can be minimal-doubling, where $\pm\beta_{j}$ correspond to locations of two nodes. The Hamiltonian (35) reduces to Borici-Creutz one (28) by choosing (29) for AA, which fixes $\beta_{j}=\pi/4$ and $\alpha=\sum_{j=1}^{3}\alpha_{j}$. The Hamiltonian (35) also yields the twisted-ordering one by choosing AA as
@@ III.1 Review of Gioia–Thorngren setup
$h^{\text{\tiny BdG}}_{\text{\tiny KW}}({\bm{p}})\,=\,\frac{1}{2}\Bigl[\sigma_{1}\,\sin p_{1}+\sigma_{3}\,\sin p_{2}+\tau_{3}\sigma_{2}\,\sin p_{3}+\sigma_{2}\,r\left(2-\cos p_{1}-\cos p_{2}\right)\Bigr],$ (41)
@@ III.3 One-parameter deformation, extra nodes and need for tuning
Because $\sigma_{1},\sigma_{2},\sigma_{3}$ anticommute and the $\tau$ matrices commute with the $\sigma$ matrices, one finds
$4h(\bm{p})^{2}=\sin^{2}p_{1}+\sin^{2}p_{2}+\Big[rM\,\mathbf{1}_{\tau}+a\tau_{3}+b\tau_{1}\Big]^{2}\,$, (59)
Thus, the eigenvalues of $4h(\bm{p})^{2}$ are
$\sin^{2}p_{1}+\sin^{2}p_{2}+\Big(rM\pm\sqrt{a^{2}+b^{2}}\Big)^{2}\,.$ (60)
@@
Thus, once the bound $|\mu|<\sqrt{r^{2}-1}$ is violated, the system is no longer in the phase with a single Weyl node: additional gapless nodes are created at $(\pi,0,\hat{p}_{3})$ and $(0,\pi,\hat{p}_{3})$.